OpenStudy (anonymous):

Find the greatest common factor for the group of terms. -66a^4,11a^5

OpenStudy (anonymous):

\[11a^4\]

OpenStudy (anonymous):

because \[11a^5=11a^4\times a\]and \[-66a^4=-6\times 11a^4\]

OpenStudy (anonymous):

I think, only 11a^4

OpenStudy (anonymous):

yes, it is \[11a^4\] my last two posts were the explanation